OpenStudy (anonymous):

I am in the middle of a question where i need to know : 2^x=-8 what would x be ????

OpenStudy (anonymous):

the answer your be 2^3 because you would mulitply 2*2*2 = 8 well thats the same thing as 2^3 which also =8

OpenStudy (anonymous):

no its -8 thats why i am stuck

jimthompson5910 (jim_thompson5910):

There are no real solutions to 2^x = -8 because 2^x is always positive for any value of x. It's impossible for 2^x to be negative (given x is a real number).

OpenStudy (anonymous):

There is no number that you can raise 2 to the power of and get a negative number. ... agree with jim.

OpenStudy (anonymous):

thats what i thought... thanks everyone !